Bluechips Inc. generates a rate of return of 18 percent on its investments and maintains a retention ratio of 0.40. Its earnings this year will be $3 per share. The required rate of return is 14 percent. a) Find the price and P/E ratio of the firm. b) What happens to the P/E ratio if the retention ratio is increased to 0.55? Why? c) Show that if the retention ratio equals zero, the earnings-price ratio, E/P, falls to the expected rate of return on the stock.
Answer #1

1.
Price=3*(1-0.4)/(14%-18%*0.4)=$26.47
P/E ratio=26.47/3=8.823333333

2.
P/E ratio=3*(1-0.55)/(14%-18%*0.55)*1/3=10.97560976
Because of higher growth of 18% through retained funds

3.
E/P ratio=1/(3*(1-0)/(14%-18%*0)*1/3)=14%
which is the expected return